Solve algebraically using one variable: The perimeter (distance around) of a rectangle is 38 inches and its area (length times width) is 70 square inches. Find the length and the width of this rectangle.

Answer by scott8148(6628)

let x equal the width ... the perimeter (2W+2L) is 38 ... so 2x+2L=38 ... 2L=38-2x ... L=19-x

the area (L*W) is 70 ... so (19-x)x=70 ... 19x-x^2=70 ... 0=x^2-19x+70 ... 0=(x-14)(x-5) ... x=14 or x=5

the two dimensions are 5 and 14 ... which is L and which is W depends on the x value you choose